OpenStudy (steve816):

The density of a substance is 570 mg/mL. What is its density in kg/L?

ganeshie8 (ganeshie8):

Hint : \(1 mg = 10^{-3}g\) \(1mL = 10^{-3}L\)

jimthompson5910 (jim_thompson5910):

it might be easier to use 1000 mg = 1 g 1000 mL = 1 L that's just my opinion though

OpenStudy (steve816):

Hmmm, I got .000570 kg/L Is that correct?

jimthompson5910 (jim_thompson5910):

This is one way to set up \[\Large \frac{570 \text{ mg}}{1 \text{ mL}} \times \frac{1000 \text{ mL}}{1 \text{ L}} \times \frac{1 \text{ g}}{1000 \text{ mg}} \]

jimthompson5910 (jim_thompson5910):

oh wait, you want kg not g, let me fix

jimthompson5910 (jim_thompson5910):

1000 mg = 1 g 1000 mL = 1 L 1000 g = 1 kg \[\Large \frac{570 \text{ mg}}{1 \text{ mL}} \times \frac{1000 \text{ mL}}{1 \text{ L}} \times \frac{1 \text{ g}}{1000 \text{ mg}} \times \frac{1 \text{ kg}}{1000 \text{ g}} \]

jimthompson5910 (jim_thompson5910):

0.000570 kg/L is not correct

OpenStudy (steve816):

I think I figured this out! Thank you jim, very helpful :)

jimthompson5910 (jim_thompson5910):

glad it's making sense now